OpenStudy (anonymous):

An electron has an initial velocity of 6.00 ✕ 106 m/s in a uniform 1.00 ✕ 105 N/C strength electric field. The field accelerates the electron in the direction opposite to its initial velocity. (b) How far does the electron travel before coming to rest? m (c) How long does it take the electron to come to rest? s (d) What is the electron's speed when it returns to its starting point? m/s

OpenStudy (anonymous):

Not sure though, \(\Delta V=Ed\\V_f-V_i=Ed\\\Large d=\frac{V_f-V_i}{E}\) where your \(V_f=0 m/s\), \(V_i=6.00 ✕ 10^6 m/s\), \(E= 1.00 ✕ 10^5 N/C\) substitute these values to solve for \(b\)

OpenStudy (anonymous):

For letter C. \(t=d/\Delta v\), where d is your answer from letter b and \(\Delta V=V_f-V_i\)

OpenStudy (schrodingers_cat):

You know F =qE =ma So, a = qE/m Then use kinematic equations Vf^2 = Vi^2 +-2ax Set Vf equal to zero and solve for x Vf = -at + Vi When Vf is zero the electron is at rest, set Vf = 0 then solve for t The object returns to its starting point when S= 0 = -1/2at^2 + vt Solve for t then substitute into Vf = - at +Vi Hope this helps :)

OpenStudy (anonymous):

You can answer the first part by considering how the kinetic and potential energies of the electron change - if you know the expressions for kinetic energy and change in potential energy in a uniform electric field. (it's quite simple)

OpenStudy (anonymous):

simple i wish

OpenStudy (anonymous):

i can do the math if i have he prper formulas

OpenStudy (anonymous):

do you remember the formula for gravitational potential energy in a uniform field, like near the surface of the earth when the accn due to gravity is g ?

OpenStudy (anonymous):

The formula is very similar in a uniform electric field.

OpenStudy (anonymous):

F G = G mM r 2

OpenStudy (anonymous):

im working on part b

OpenStudy (anonymous):

using d= vf-vi/e but its not correct answer

OpenStudy (anonymous):

ok - never mind the gravity thing how do you work out the acceleration of the electron from the numbers given in the question ?

OpenStudy (anonymous):

velocity/distance

OpenStudy (anonymous):

v/e ?

OpenStudy (anonymous):

what is the force on a charge q in an electric field E ?

OpenStudy (anonymous):

coulombs law

OpenStudy (anonymous):

no, coulombs law gives the force between 2 charges, but we are looking at a single charge in a field E

OpenStudy (anonymous):

you have to know what force a charge q feels in a field E to answer this question

OpenStudy (anonymous):

its a much simpler expression than coulombs law

OpenStudy (anonymous):

e=f/q

OpenStudy (anonymous):

well yes, i would have said f = qE, same thing

OpenStudy (anonymous):

f=qe

OpenStudy (anonymous):

ok, now if a mass m experiences a force f, what is its acceleration, in terms of f and m ?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

newtons second law tells you

OpenStudy (anonymous):

a=f/m o yea

OpenStudy (anonymous):

right !

OpenStudy (anonymous):

and in our problem, we know that the force f is qE, like we just discussed above

OpenStudy (anonymous):

so, you can put those pieces together and find that a =qE/m for an electron of charge q and mass m in an electric field E

OpenStudy (anonymous):

Now you're in business, because that a can be used in your formulae for motion in a straight line under constant acceleration

OpenStudy (anonymous):

\[v ^{2}=u ^{2}+2as\] and so on

OpenStudy (anonymous):

but how do i get distance dng im supid

OpenStudy (anonymous):

well what do we know ? we know the velocity at the start of the motion and at the end of the distance covered

OpenStudy (anonymous):

u is given in the question, and we want v = 0, because that's what 'coming to rest' means

OpenStudy (anonymous):

i got that

OpenStudy (anonymous):

ok and you can calculate a from the numbers given in the question

OpenStudy (anonymous):

so that only leaves s to be worked out from the equation i just put up

OpenStudy (anonymous):

whats the s represent?

OpenStudy (anonymous):

the distance !

OpenStudy (anonymous):

im used to r or d o

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

: ) it's ok

OpenStudy (anonymous):

so i use a=f/m to find find aaccelleration mbieng the mass of an electron?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and re work the v^2=u^2+as to find the difference

OpenStudy (anonymous):

v= velocity u = uniform field and a =cceloration and s bieng distance

OpenStudy (anonymous):

u is initial velocity v is final velocity

OpenStudy (anonymous):

aww so i dont use the 1e5 of the field no where?

OpenStudy (anonymous):

you need that to calculate f

OpenStudy (anonymous):

f being the force, f = Eq, E being the field

OpenStudy (anonymous):

whch is udes to find a

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

o starting to click thanks a bunch

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

q bieng the 1.6e-19

OpenStudy (anonymous):

yes, that's right, the charge on the electron

OpenStudy (anonymous):

e being the 1E5

OpenStudy (anonymous):

usually written E for electric field, but yes

OpenStudy (anonymous):

tanks again your a genuis and thanks forhelping me understand instead of here your answer

OpenStudy (anonymous):

you're welcome, i hope it starts to seem clearer for you - btw i'm nowhere on the genius scale - the needle on the genius detector wouldn't even flicker

OpenStudy (anonymous):

i seeing the relationships any way lol

OpenStudy (anonymous):

good

OpenStudy (anonymous):

well u know the lectic aspect very well

OpenStudy (anonymous):

so inorder to rearrange the v2=u2+as it would be 0=6e6^2+1.76e16*?

OpenStudy (anonymous):

find s

OpenStudy (anonymous):

6e6^2/1.76216 ?

OpenStudy (anonymous):

still here?

OpenStudy (anonymous):

ok, tell me what you have for a, the electron acceleration, the numerical value

OpenStudy (anonymous):

1.756e16

OpenStudy (anonymous):

that is correct so the other number you need is u squared , what do you have for that ?

OpenStudy (anonymous):

3.6e13

OpenStudy (anonymous):

then divide

OpenStudy (anonymous):

also correct so now , ignoring the minus sign, we have s = u squared/2a you've got the right numbers so far, so if you plug them in you should get the right value for s

OpenStudy (anonymous):

don't forget the factor of 2 from the equation

OpenStudy (anonymous):

haha crap i didnt write that down earlier dang it

OpenStudy (anonymous):

the 2 tat is

OpenStudy (anonymous):

thats what got me du once again you da boom

OpenStudy (anonymous):

it should come out at about 1e-3 meters, about a millimetre - does that look about right to you ?

OpenStudy (anonymous):

i get 3.16e29

OpenStudy (anonymous):

what !!!!!!!!!!! you must have miscalculated lol

OpenStudy (anonymous):

3.6e13/(2 x 1.76e16)

OpenStudy (anonymous):

3.16e13/2(1.765e16)

OpenStudy (anonymous):

its roughly 1e-3

OpenStudy (anonymous):

opps parent in wrong order

OpenStudy (anonymous):

did it wrong on the the ol t-83

OpenStudy (anonymous):

yes i got that

OpenStudy (anonymous):

the next part you can get easily cos you know u, v and a, there's another formula that will give you the time from those three

OpenStudy (anonymous):

and for the last part there's a trick you can use - you don't need to calculate anything

OpenStudy (anonymous):

for partc t=d/v

OpenStudy (anonymous):

that would only work for constant speed how about using v=u+at

OpenStudy (anonymous):

and the trick for last part

OpenStudy (anonymous):

i'll give you a hint - it's about the total energy of the electron

OpenStudy (anonymous):

e=mc^2 ?

OpenStudy (anonymous):

its like the first problem except we find final instead of intal

OpenStudy (anonymous):

the total energy is conserved

OpenStudy (anonymous):

right isnt lost

OpenStudy (anonymous):

so when the electron comes back to its starting point, but moving in the opposite direction, at that point it has the same kinetic energy it had when it set off

OpenStudy (anonymous):

so same as vi